8) As part of a study of corporate employees, the Director of Human Resources for QRS Company wants to compare the distance traveled to work by employees at their office in downtown Cincinnati with the distance for those in downtown Pittsburgh. a) State the Hypothesis to show mean number of miles traveled per month is different for Cincinnati and Pittsburgh employees. b) Choose a level of a. Use a = 0.05 for this problem. c) To test the hypothesis, the researcher collects data on 35 workers in the Cincinnati office and 40 in the Pittsburg office. The data appear in the QRS Company worksheet of the HW1 data workbook on Moodle. Collect data and calculate necessary statistics to test the hypothesis. d) Sketch the sampling distribution. Include the critical value and test statistic e) Draw a conclusion and report that in the problem context. f) Calculate the p-value for the hypothesis test g) Use Excel to perform the hypothesis test.

Cincinnati Pittsburg
406 347
333 379
346 360
331 360
400 379
368 339
350 385
384 353
351 397
369 369
368 418
362 367
334 391
313 387
322 421
335 370
412 353
317 353
378 376
415 382
376 410
434 371
388 402
408 402
377 354
371 391
382 412
367 305
365 385
363 412
410 346
367 411
374 401
407 393
367 360
366
409
389
372
423
Answer #1

a) NULL HYPOTHESIS H0: mu (C)= mu (P)

ALTERNATIVE HYPOTHESIS Ha: mu (C) e mu (P)

b) Level of Significance (alpha)= 0.05

c) To test the hypothesis we will assume following assumptions:

i) The samples are independent to each other.

ii) The samples are normally distributed.

iii) Populations Variances are unknown and equal to each other.

To verify last assumption we will use F test .

H0: POPULATION VARIANCES ARE EQUAL

Ha: POPULATION VARIANCES ARE NOT EQUAL.

F-Test Two-Sample for Variances CINNCINATI PITTSBURG Mean Variance Observations df 370 900.2352941 35 34 1.330901307 0.193647721 1.728029639 380 676.4102564 40 39 P(F< f) one-tail F Critical one-tail

From above test we can see that P value=0.1936 >0.05 therefore NOT SIGNIFICANT.

We can conclude that population variances are equal.

Now

t-Test: Two-Sample Assuming Equal Variances Cinncinati Pittsbur Mean Variance Observations Pooled Variance Hypothesized Mean D df t Stat PlTct) one-tail t Critical one-tail P(T<-t) two-tail t Critical two-tail 370 900.2352941 35 780.6575342 0 73 1.5463333 0.063174292 1.665996224 0.126348584 1.992997097 380 676.4102564 40

From above table we can see that

t stat= -1.546

t stat= 1.546 (Two tailed test)

t critical= 1.99

P value =0.1263

Since P value >0.05 level of significance therefore NOT significant.

Decision: Fail to reject H0.

Conclusion: We don't have enough evidence to conclude that mean number of miles traveled per month is different for Cincinnati and Pittsburgh employees.

P-Value=0. I 264 -5 -3 .2 -1 0 test t=1.546 t-1.99 t= 1 .99